# Finite-Agent Stochastic Differential Games on Large Graphs: I. The Linear-Quadratic Case

### v1

@@ 3.1 The mathematical formulation
Therefore, the FP procedure for the game introduced in Section 2.1 has been completely described by system (3.9) following from Assumption 3.2.
@@ 3.2 Convergence results
In this section, we prove that the iterative scheme described by (3.9) converges to a limit, which is the Markovian NE. We first define a notion of convergence when discussing the convergence of strategies. Since the players’ strategies are represented by feedback functions of state processes, we examine the pointwise convergence of the feedback function $\phi^{i,k}(t,x)$ as $k\to\infty$ for any fixed $i\in[N]$. As implied by Lemma 3.3 and equation (3.8), the pointwise convergence of $\phi^{i,k}$ is equivalent to the convergence of $F^{i,k}$. Therefore, it suffices to prove that the solution $F^{i,k}$ to system (3.9) admits a limit as $k\to\infty$ for any fixed $i\in[N]$ (Theorem 3.6), and the limit satisfies equation (2.13) (Theorem 3.7). The proof relies on the following two technical lemmas. The first lemma provides a way to prove the uniform convergence of a function sequence.
@@
Let $f_{k}:[0,T]\to\mathbb{R}_{+}$ be a sequence of continuous non-negative functions. If there exists a constant $C>0$ such that
$f_{k}(t)\leq C\int_{t}^{T}[f_{k}(s)+f_{k-1}(s)]\,\mathrm{d}s,\ \forall t\in[0,T],\ \forall k\in\mathbb{N},$ (3.10)
then $f_{k}$ converges to $f\equiv 0$, uniformly in $t\in[0,T]$, as $k\to\infty$.
Proof.
Consider $g_{k}(t):=e^{\beta t}f_{k}(t)$ for any $\beta>0$. It is clear that
$g_{k}(t)\leq Ce^{\beta t}\int_{t}^{T}e^{-\beta s}[g_{k}(s)+g_{k-1}(s)]\,\mathrm{d}s.$ (3.11)
Since $g_{k}(t)$ is a continuous function on a compact domain, $\sup_{s\in[0,T]}g_{k}(s)<\infty$, which provides an upper bound for the integral above:
$g_{k}(t)\leq\frac{C}{\beta}(1-e^{-\beta(T-t)})\left(\sup_{s\in[0,T]}g_{k}(s)+\sup_{s\in[0,T]}g_{k-1}(s)\right).$ (3.12)
Specifying the value of $\beta$ to be large enough such that $\frac{C}{\beta}<\frac{1}{3}$, and taking the supremum on both sides with respect to $t\in[0,T]$ yields
$\sup_{t\in[0,T]}g_{k}(t)\leq\frac{1}{2}\sup_{t\in[0,T]}g_{k-1}(t),$ (3.13)
which implies that
$\sup_{t\in[0,T]}g_{k}(t)\leq\frac{1}{2^{k}}\sup_{t\in[0,T]}g_{0}(t)\to 0\ (k\to\infty).$ (3.14)
Note that $\sup_{t\in[0,T]}f_{k}(t)\leq\sup_{t\in[0,T]}g_{k}(t)$, setting $k\to\infty$ concludes the proof. ∎
The second lemma proves a matrix inequality. Unless otherwise specified, all matrix norms are understood as the matrix $2$-norm, and the inequality between symmetric matrices is defined in terms of the semi-positive-definite sense. That is, $A\geq B$ if and only if $A-B$ is semi-positive-definite, for $A,B\in\mathbb{S}^{N\times N}$.
Lemma 3.5.
For symmetric semi-positive-definite matrices $A_{1},\ldots,A_{N}$, the following inequality holds:
@@
holds, then $\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|$ converges to 00, uniformly in $t\in[0,T]$, as $k\to\infty$, where $F^{i,k}_{t}$ is the solution to system (3.9).
Consequently, $F_{t}^{i,k}$ converges for any fixed $i\in[N]$, uniformly in $t\in[0,T]$, as $k\to\infty$. We denote the uniform limit as $F^{i,\infty}_{t}:=\lim_{k\to\infty}F^{i,k}_{t}$.
@@
$\dot{W}^{i,k+1}_{t}=-(W^{i,k+1}_{t}-W^{i,k}_{t})e_{i}e_{i}^{\operatorname{T}}(W^{i,k+1}_{t}-W^{i,k}_{t})+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-(a+q)LW^{i,k+1}_{t}-(a+q)W^{i,k+1}_{t}L\\ -W^{i,k+1}_{t}\sum_{j=1}^{N}e_{j}e_{j}^{\operatorname{T}}W^{j,k}_{t}-\sum_{j=1}^{N}W^{j,k}_{t}e_{j}e_{j}^{\operatorname{T}}W^{i,k+1}_{t}+W^{i,k}_{t}e_{i}e_{i}^{\operatorname{T}}W^{i,k}_{t},\quad W^{i,k+1}_{0}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (3.19)
Using $W^{i,k+1}_{t}=W^{i,k+1}_{0}+\int_{0}^{t}\dot{W}^{i,k+1}_{s}\,\mathrm{d}s$ and $(W^{i,k+1}_{t}-W^{i,k}_{t})e_{i}e_{i}^{\operatorname{T}}(W^{i,k+1}_{t}-W^{i,k}_{t})\geq 0$, followed by summing both sides with respect to $i\in[N]$, gives the following matrix inequality:
@@
Let $A^{k}_{t}:=\sum_{i=1}^{N}W^{i,k}_{t}$ and $B^{k}_{t}:=\sum_{i=1}^{N}e_{i}e_{i}^{\operatorname{T}}W^{i,k}_{t}$. The matrix inequality above is rewritten as
@@
For any $i\in[N]$, $k\in\mathbb{N}$, the cost functionals $f^{i}$ and $g^{i}$ stay non-negative, so does the value function $v^{i,k}(t,x)=\frac{1}{2}x^{\operatorname{T}}F^{i,k}_{t}x+h^{i,k}_{t}$, $\forall x\in\mathbb{R}^{N}$. Consequently, one has $F^{i,k}_{t}\geq 0$, $W^{i,k}_{t}\geq 0$, $\forall t\in[0,T]$. Then Lemma 3.15 tells $\|B^{k}_{t}\|\leq\|A^{k}_{t}\|$. Combining with inequality (3.21), we have
@@
$\sup_{t\in[0,T]}\|A^{k+1}_{t}\|\leq\left([c+T(\varepsilon-q^{2})]\;\|L\|^{2}+T\sup_{t\in[0,T]}\|A^{k}_{t}\|^{2}\right)e^{2T(a+q)\|L\|+2T\sup_{t\in[0,T]}\|A^{k}_{t}\|}.$ (3.24)
@@
Under condition (3.18), $[c+T(\varepsilon-q^{2})]\;\|L\|^{2}\leq 1$ and $e^{2T(a+q)\|L\|}\leq 2$, so
@@
Since $g(2)>2$, $g(3)<3$, by the intermediate value theorem, $g$ has a fixed point $x^{*}\in(2,3)$. Since FP starts from $F^{i,0}\equiv 0$, $\sup_{t\in[0,T]}\|A^{0}_{t}\|=0$. A simple induction argument implies that
$\sup_{t\in[0,T]}\|A^{k+1}_{t}\|\leq h\left(\sup_{t\in[0,T]}\|A^{k}_{t}\|\right)\leq g\left(\sup_{t\in[0,T]}\|A^{k}_{t}\|\right)\leq g(x^{*})=x^{*},\ \forall k\in\mathbb{N}.$ (3.27)
This inequality is equivalent to $\|\sum_{i=1}^{N}F^{i,k}_{t}\|\leq x^{*}$, $\forall t\in[0,T]$, $\forall k\in\mathbb{N}$. Combining with the fact that $F^{i,k}_{t}$ takes values as semi-positive-definite matrices proves the a priori bound $\|F^{i,k}_{t}\|\leq x^{*}$, $\forall k\in\mathbb{N}$, $\forall t\in[0,T]$, $\forall i\in[N]$.
@@
$\|\Delta F^{i,k}_{t}\|\leq 2(a+q)\|L\|\int_{t}^{T}\|\Delta F^{i,k}_{s}\|\,\mathrm{d}s+\int_{t}^{T}\left(\|F^{i,k}_{s}\|\cdot\|\Delta F^{i,k}_{s}\|+\|F^{i,k+1}_{s}\|\cdot\|\Delta F^{i,k}_{s}\|\right)\,\mathrm{d}s\\ +2\int_{t}^{T}\left(\|\Delta F^{i,k}_{s}\|\cdot\sum_{j\neq i}\|F^{j,k-1}_{s}\|+\|F^{i,k+1}_{s}\|\cdot\sum_{j\neq i}\|\Delta F^{j,k-1}_{s}\|\right)\,\mathrm{d}s.$ (3.28)
@@
$\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|\leq\left[2(a+q)\|L\|+2x^{*}+2Nx^{*}\right]\int_{t}^{T}\sup_{i\in[N]}\|\Delta F^{i,k}_{s}\|\,\mathrm{d}s+2Nx^{*}\int_{t}^{T}\sup_{i\in[N]}\|\Delta F^{i,k-1}_{s}\|\,\mathrm{d}s.$ (3.29)
By Lemma 3.4, $\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|$ converges to 00, uniformly in $t\in[0,T]$, as $k\to\infty$. This concludes the proof. ∎
In addition, we argue that the limit $F_{t}^{i,\infty}$ identified in Theorem 3.6 corresponds to the Markovian NE, formulated below as Theorem 3.7.
Theorem 3.7.
Under Assumption 3.2, if the fictitious play starts from $F^{i,0}\equiv 0$ for any $i\in[N]$ and condition (3.18) holds, then the scheme (3.9) produced by fictitious play converges to the Markovian NE for the linear-quadratic games on graphs (2.2)–(2.4).
Proof.
It suffices to prove that $(F^{1,\infty}_{t},\cdots,F^{N,\infty}_{t})$, the limit of $(F^{1,k}_{t},\cdots,F^{N,k}_{t})$ as $k\to\infty$, is a solution to the coupled Riccati system (2.13). Set $k\to\infty$ in equation (3.9), and it remains to justify the interchange of the differentiation with respect to $t$ and the limit with respect to $k$. To this end, we need to show $\lim_{k\to\infty}\dot{F}^{i,k}_{t}=\dot{F}^{i,\infty}_{t}$.
Rewrite equation (3.9) as $\dot{F}^{i,k+1}_{t}=H(F^{i,k+1}_{t},F^{1,k}_{t},\cdots,F^{i-1,k}_{t},F^{i+1,k}_{t},\cdots,F^{N,k}_{t})$ for some continuous function $H:(\mathbb{S}^{N\times N})^{N}\to\mathbb{S}^{N\times N}$. From the proof of Theorem 3.6, one has $\|F^{i,k}_{t}\|\leq x^{*}$ for every $k\in\mathbb{N}$, $i\in[N]$, $t\in[0,T]$. Hence, the domain of $H$ is compact, and $H$ is uniformly continuous. From Theorem 3.6, the convergence of $F^{i,k}_{t}$ is uniform w.r.t. $t\in[0,T]$. It’s well-known that uniformly continuous transformations preserve uniform convergence. As a result, $\dot{F}^{i,k+1}_{t}$ converges uniformly w.r.t. $t\in[0,T]$ as $k\to\infty$. The interchange of the limit and the differentiation is justified, which concludes the proof. ∎
@@
$\dot{G}^{i,k+1}_{t}-G^{i,k+1}_{t}e_{i}e_{i}^{\operatorname{T}}G^{i,k+1}_{t}+\left(\frac{T}{T^{\prime}}\right)^{2}(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-\frac{T}{T^{\prime}}(a+q)LG^{i,k+1}_{t}-\frac{T}{T^{\prime}}(a+q)G^{i,k+1}_{t}L\\ -G^{i,k+1}_{t}\sum_{j\neq i}e_{j}e_{j}^{\operatorname{T}}G^{j,k}_{t}-\sum_{j\neq i}G^{j,k}_{t}e_{j}e_{j}^{\operatorname{T}}G^{i,k+1}_{t}=0,\quad G^{i,k+1}_{T^{\prime}}=\frac{T}{T^{\prime}}cLe_{i}e_{i}^{\operatorname{T}}L.$ (3.31)
This implies that the FP procedure for the linear-quadratic games on graphs (2.2)–(2.4) with model parameters $(T,a,q,\varepsilon,c)$ is equivalent to the one with model parameters $(T^{\prime},a^{\prime},q^{\prime},\varepsilon^{\prime},c^{\prime})$, where
@@
The condition (3.30) imposes a smallness constraint on the model parameters, a requirement that is commonly found in the literature. For instance, the convergence of FP to the open-loop NE for linear-quadratic games on complete graphs requires that $T$ or $(c,q,\epsilon-q^{2})$ are sufficiently small, or $a$ is sufficiently large [Hu21]. The convergence of the Deep BSDE method [EHJ17] for stochastic control problems with Lipschitz models requires a small time duration or small Lipschitz coefficients [HHL22], as well as the well-posedness of forward-backward stochastic differential equations (FBSDEs) [MWZZ15]. Condition (3.30) aligns with these requirements, and it is certainly not the strictest constraint and can be easily extended. As long as reasonable bounds are imposed on model parameters such that the function $h$ in equation (3.25) has a fixed point in $\mathbb{R}_{+}$, Theorems 3.63.7 and Corollary 3.8 remain valid and provide a standard way to find a condition, under which FP is guaranteed to converge. Lastly, we want to emphasize two points: (i) FP does not necessarily have to start from $F^{i,0}\equiv 0$. From the proof of Theorem 3.6, the convergence results still hold as long as $\sup_{t\in[0,T]}\|\sum_{i=1}^{N}F^{i,0}_{t}\|\leq x^{*}$. (ii) Condition (3.30) does not impose any restrictions on the number of players $N$. The convergence result still holds even when there is a large number of players, i.e., $N\to\infty$.
@@
We are particularly interested in the connection between the graph structure and the convergence of fictitious play (FP). This relationship can be analyzed through $C_{L}$ defined in condition (3.30). Fixing the other model parameters, an increase in $\|L\|$ leads to an increase in $C_{L}$, making it more likely to violate condition (3.30). Taking a closer look, we note that any graph Laplacian $L$ must satisfy $0\leq\|L\|\leq 2$. If a graph has no isolated vertices, its graph Laplacian satisfies $\|L\|\geq\frac{N}{N-1}$ [Chu97]. A connected bipartite graph has the norm of the graph Laplacian achieve the upper bound $\|L\|=2$, while a complete graph achieves the lower bound $\|L\|=\frac{N}{N-1}$. Although the convergence rate of FP has not been accurately quantified in Corollary 3.8, numerical experiments show a connection between the graph structure and the convergence rate of FP. We refer the readers to Section 5.2 for a more detailed discussion.

### v2

@@ 3.1 The mathematical formulation
Therefore, the FP procedure for the game introduced in Section 2.1 has been completely described by system (3.9) following from Assumption 3.3. In Proposition 3.5 we prove that, given $\{F^{j,k}\}_{j\in[N]}$, system (3.9) uniquely determines $\{F^{j,k+1}\}_{j\in[N]}$ on $[0,T]$ for any $T>0$, ensuring that the FP iteration scheme is well-defined.
Unless otherwise specified, all matrix norms are understood as the matrix $2$-norm, and inequalities between symmetric matrices are understood in the positive semi-definite sense. That is, for $A,B\in\mathbb{S}^{N\times N}$, we write $A\geq B$ if and only if ABA-B is positive semi-definite.
Proposition 3.5.
Given $\{F^{j,0}\}_{j\in[N]}$ such that $F^{j,0}_{t}$ is bounded on $[0,T]$ for any $j\in[N]$, system (3.9) admits a unique solution $\{F^{j,k+1}\}_{j\in[N]}$ on $[0,T]$ for any $T>0$ and $k\in\mathbb{N}$. Moreover, $F^{j,k+1}_{t}\geq 0$ for any $t\in[0,T]$, $k\in\mathbb{N}$, $j\in[N]$.
Proof.
We prove the result by induction on $k$. Suppose that $\{F^{j,k}\}_{j\in[N]}$ is given. Define the function $f:[0,T]\times\mathbb{S}^{N\times N}\to\mathbb{S}^{N\times N}$ such that
$f(t,X;\{F^{j,k}\}_{j\in[N]}):=Xe_{i}e_{i}^{\operatorname{T}}X-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L+(a+q)LX+(a+q)XL\\ +X\sum_{j\neq i}e_{j}e_{j}^{\operatorname{T}}F^{j,k}_{t}+\sum_{j\neq i}F^{j,k}_{t}e_{j}e_{j}^{\operatorname{T}}X.$ (3.10)
Then system (3.9) can be rewritten as $\dot{F}^{i,k+1}_{t}=f(t,F^{i,k+1}_{t};\{F^{j,k}\}_{j\in[N]})$ with the terminal condition $F^{i,k+1}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L$.
We claim that: $f$ is Lipschitz continuous in $X$ if there exists $M>0$ such that $\|X\|\leq M$. This follows directly from the triangle inequality:
$\|f(t,X_{1})-f(t,X_{2})\|\leq 2\left(M+(a+q)\|L\|+\|\sum_{j\neq i}e_{j}e_{j}^{\operatorname{T}}F^{j,k}_{t}\|\right)\|X_{1}-X_{2}\|.$ (3.11)
Since $f$ is continuous in $t$, the Picard-Lindelöf theorem [Tes12] ensures that system (3.9) has a unique local solution at time $T$.
To achieve the global existence and uniqueness result, it suffices to establish the a priori boundedness of $F^{i,k+1}_{t}$. Consider the auxiliary ODE system for $\{G^{j,k+1}\}_{j\in[N]}$:
$\dot{G}^{i,k+1}_{t}+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-(a+q)LG^{i,k+1}_{t}-(a+q)G^{i,k+1}_{t}L\\ -G^{i,k+1}_{t}\sum_{j\neq i}e_{j}e_{j}^{\operatorname{T}}F^{j,k}_{t}-\sum_{j\neq i}F^{j,k}_{t}e_{j}e_{j}^{\operatorname{T}}G^{i,k+1}_{t}=0,\quad G^{i,k+1}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (3.12)
Since this is a linear ODE system with bounded time-variant coefficients, it admits a unique bounded solution $\{G^{j,k+1}\}_{j\in[N]}$ on $[0,T]$ for any $T>0$. Consider the difference $D^{i,k+1}_{t}:=F^{i,k+1}_{t}-G^{i,k+1}_{t}$, which satisfies $D^{i,k+1}_{T}=0$ and $\dot{D}^{i,k+1}_{t}\geq 0$. Therefore, $D^{i,k+1}_{t}\leq 0$, implying an upper bound $F^{i,k+1}_{t}\leq G^{i,k+1}_{t}$, $\forall t\in[0,T]$.
For any $j\in[N]$, $k\in\mathbb{N}$, the cost functionals $f^{j}$ and $g^{j}$ stay non-negative, so does the value function $v^{j,k+1}(t,x)=\frac{1}{2}x^{\operatorname{T}}F^{j,k+1}_{t}x+h^{j,k+1}_{t}$, $\forall x\in\mathbb{R}^{N}$. Consequently, one obtains the lower bound $F^{j,k+1}_{t}\geq 0$, $\forall t\in[0,T]$.
Combining both parts and setting $M:=\sup_{t\in[0,T]}\sup_{j\in[N]}\|G^{j,k+1}_{t}\|$ conclude the proof. ∎
@@ 3.2 Convergence results
In this section, we prove that the iterative scheme described by (3.9) converges to a limit, which is the Markovian NE. We first define a notion of convergence when discussing the convergence of strategies. Since the players’ strategies are represented by feedback functions of state processes, we examine the pointwise convergence of the feedback function $\phi^{i,k}(t,x)$ as $k\to\infty$ for any fixed $i\in[N]$. As implied by Lemma 3.4 and equation (3.8), the pointwise convergence of $\phi^{i,k}$ is equivalent to the convergence of $F^{i,k}$. Therefore, it suffices to prove (in Theorem 3.7) that the solution $F^{i,k}$ to system (3.9) admits a limit as $k\to\infty$ for any fixed $i\in[N]$ , and the limit satisfies equation (2.14). The proof relies on the following technical lemma as a matrix inequality.
@@
For symmetric positive semi-definite matrices $A_{1},\ldots,A_{N}$, the following inequality holds:
@@
holds, then the scheme (3.9) produced by fictitious play converges to a Markovian NE for the linear-quadratic games on graphs (2.2)–(2.4).
@@
$\dot{W}^{i,k+1}_{t}=-(W^{i,k+1}_{t}-W^{i,k}_{t})e_{i}e_{i}^{\operatorname{T}}(W^{i,k+1}_{t}-W^{i,k}_{t})+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-(a+q)LW^{i,k+1}_{t}\\ -(a+q)W^{i,k+1}_{t}L-W^{i,k+1}_{t}\sum_{j=1}^{N}e_{j}e_{j}^{\operatorname{T}}W^{j,k}_{t}-\sum_{j=1}^{N}W^{j,k}_{t}e_{j}e_{j}^{\operatorname{T}}W^{i,k+1}_{t}+W^{i,k}_{t}e_{i}e_{i}^{\operatorname{T}}W^{i,k}_{t},$ (3.17)
with terminal condition $W^{i,k+1}_{0}=cLe_{i}e_{i}^{\operatorname{T}}L$. Using $W^{i,k+1}_{t}=W^{i,k+1}_{0}+\int_{0}^{t}\dot{W}^{i,k+1}_{s}\,\mathrm{d}s$ and $(W^{i,k+1}_{t}-W^{i,k}_{t})e_{i}e_{i}^{\operatorname{T}}(W^{i,k+1}_{t}-W^{i,k}_{t})\geq 0$, followed by summing both sides with respect to $i\in[N]$, gives the following matrix inequality:
@@
Let $A^{k}_{t}:=\sum_{j=1}^{N}W^{j,k}_{t}$ and $B^{k}_{t}:=\sum_{j=1}^{N}e_{j}e_{j}^{\operatorname{T}}W^{j,k}_{t}$. The matrix inequality above is rewritten as
@@
Since Proposition 3.5 implies that $W^{i,k}_{t}\geq 0$, $\forall t\in[0,T]$, Lemma 3.13 tells $\|B^{k}_{t}\|\leq\|A^{k}_{t}\|$. Combining with inequality (3.19), we have
@@
$\sup_{t\in[0,T]}\|A^{k+1}_{t}\|\leq\bigg{(}[c+T(\varepsilon-q^{2})]\;\|L\|^{2}+T\sup_{t\in[0,T]}\|A^{k}_{t}\|^{2}\bigg{)}e^{2T(a+q)\|L\|+2T\sup_{t\in[0,T]}\|A^{k}_{t}\|}.$ (3.22)
@@
Under condition (3.16), $c+T(\varepsilon-q^{2})]\;\|L\|^{2}\leq 1$ and $e^{2T(a+q)\|L\|}\leq 2$, so
@@
Since $g(2)>2$, $g(3)<3$, by the intermediate value theorem, $g$ has a fixed point $x^{*}\in(2,3)$.
Using the fixed point $x^{*}$, we establish a uniform a priori bound for $\|A^{k}_{t}\|$. Specifically, we claim that $\sup_{t\in[0,T]}\|A^{k}_{t}\|\leq x^{*},\ \forall k\in\mathbb{N}$, which follows from straightforward induction argument. For the base case $k=0$, since FP starts from $F^{i,0}\equiv 0$, we have $\sup_{t\in[0,T]}\|A^{0}_{t}\|=0\leq x^{*}$ that confirms the claim. Now, assume the claim holds for $k=l$, i.e., $\sup_{t\in[0,T]}\|A^{l}_{t}\|\leq x^{*}$. By applying the recursive structure, we obtain
$\sup_{t\in[0,T]}\|A^{l+1}_{t}\|\leq h\left(\sup_{t\in[0,T]}\|A^{l}_{t}\|\right)\leq g\left(\sup_{t\in[0,T]}\|A^{l}_{t}\|\right)\leq g(x^{*})=x^{*},$ (3.25)
where the final inequality follows from the induction hypothesis $\sup_{t\in[0,T]}\|A^{l}_{t}\|\leq x^{*}$ and the monotonicity of $g$. Consequently, we conclude that $\|\sum_{j=1}^{N}F^{j,k}_{t}\|\leq x^{*}$, $\forall t\in[0,T]$, $\forall k\in\mathbb{N}$. Combining with the fact that $F^{i,k}_{t}$ takes values as positive semi-definite matrices proves the a priori bound $\|F^{j,k}_{t}\|\leq x^{*}$, $\forall k\in\mathbb{N}$, $\forall t\in[0,T]$, $\forall j\in[N]$.
@@
$\|\Delta F^{i,k}_{t}\|\hskip-2.0pt\leq 2(a+q)\|L\|\int_{t}^{T}\|\Delta F^{i,k}_{s}\|\,\mathrm{d}s+\int_{t}^{T}\left(\|F^{i,k}_{s}\|\|\Delta F^{i,k}_{s}\|+\|F^{i,k+1}_{s}\|\|\Delta F^{i,k}_{s}\|\right)\,\mathrm{d}s\\ +2\int_{t}^{T}\left(\|\Delta F^{i,k}_{s}\|\sum_{j\neq i}\|F^{j,k-1}_{s}\|+\|F^{i,k+1}_{s}\|\sum_{j\neq i}\|\Delta F^{j,k-1}_{s}\|\right)\,\mathrm{d}s.$ (3.26)
@@
$\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|\leq\left[2(a+q)\|L\|+2x^{*}+2Nx^{*}\right]\int_{t}^{T}\sup_{i\in[N]}\|\Delta F^{i,k}_{s}\|\,\mathrm{d}s\\ +2Nx^{*}\int_{t}^{T}\sup_{i\in[N]}\|\Delta F^{i,k-1}_{s}\|\,\mathrm{d}s.$ (3.27)
Define $g_{k}(t):=e^{\beta t}\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|$ for some sufficiently large $\beta>0$ such that $\frac{2(a+q)\|L\|+2x^{*}+2Nx^{*}}{\beta}<\frac{1}{3}$. It follows that
$g_{k}(t)\leq[2(a+q)\|L\|+2x^{*}+2Nx^{*}]e^{\beta t}\int_{t}^{T}e^{-\beta s}[g_{k}(s)+g_{k-1}(s)]\,\mathrm{d}s.$ (3.28)
Since $g_{k}(t)$ is a continuous function on a compact domain, it satisfies $\sup_{s\in[0,T]}g_{k}(s)<\infty$, which provides an upper bound for the integral above:
$g_{k}(t)\leq\frac{2(a+q)\|L\|+2x^{*}+2Nx^{*}}{\beta}(1-e^{-\beta(T-t)})\bigg{(}\sup_{s\in[0,T]}g_{k}(s)+\sup_{s\in[0,T]}g_{k-1}(s)\bigg{)}.$ (3.29)
Taking the supremum with respect to $t\in[0,T]$ on both sides yields the contraction
$\sup_{t\in[0,T]}g_{k}(t)\leq\frac{1}{2}\sup_{t\in[0,T]}g_{k-1}(t).$ (3.30)
This implies that $\Delta F^{k}$ converges to zero under the norm $\|\Delta F^{k}\|_{\beta}:=\sup_{t\in[0,T]}e^{\beta t}\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|$ as $k\to\infty$. Since, on a finite time horizon $[0,T]$, the norm $\|\cdot\|_{\beta}$ is equivalent to $\|\cdot\|_{0}$, where $\|\Delta F^{k}\|_{0}:=\sup_{t\in[0,T]}\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|$, this implies that $F_{t}^{i,k}$ converges uniformly in $t\in[0,T]$, for any fixed $i\in[N]$, as $k\to\infty$. We denote the uniform limit by $F^{i,\infty}_{t}:=\lim_{k\to\infty}F^{i,k}_{t}$.
[…]
Rewriting equation (3.9), we express
$\dot{F}^{i,k+1}_{t}=H(F^{i,k+1}_{t},F^{1,k}_{t},\cdots,F^{i-1,k}_{t},F^{i+1,k}_{t},\cdots,F^{N,k}_{t}),$
where $H:(\mathbb{S}^{N\times N})^{N}\to\mathbb{S}^{N\times N}$ is a continuous function explicitly determined by the structure of equation (3.9). Since $\|F^{j,k}_{t}\|\leq x^{*}$ for all $k\in\mathbb{N}$, $j\in[N]$, and $t\in[0,T]$, the domain of $H$ is compact, and $H$ is uniformly continuous. It is well-known that uniformly continuous transformations preserve uniform convergence. Since the convergence of $F^{i,k}_{t}$ is uniform in $t\in[0,T]$, it follows that $\dot{F}^{i,k+1}_{t}$ also converges uniformly in $t\in[0,T]$ as $k\to\infty$. This justifies the interchange of the limit and differentiation, completing the proof. ∎
@@
$\dot{G}^{i,k+1}_{t}-G^{i,k+1}_{t}e_{i}e_{i}^{\operatorname{T}}G^{i,k+1}_{t}+\left(\frac{T}{T^{\prime}}\right)^{2}(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-\frac{T}{T^{\prime}}(a+q)LG^{i,k+1}_{t}\\ -\frac{T}{T^{\prime}}(a+q)G^{i,k+1}_{t}L-G^{i,k+1}_{t}\sum_{j\neq i}e_{j}e_{j}^{\operatorname{T}}G^{j,k}_{t}-\sum_{j\neq i}G^{j,k}_{t}e_{j}e_{j}^{\operatorname{T}}G^{i,k+1}_{t}=0,$ (3.32)
with terminal condition $G^{i,k+1}_{T^{\prime}}=\frac{T}{T^{\prime}}cLe_{i}e_{i}^{\operatorname{T}}L$. This implies that the FP procedure for the linear-quadratic games on graphs (2.2)–(2.4) with model parameters $(T,a,q,\varepsilon,c)$ is equivalent to the one with model parameters $(T^{\prime},a^{\prime},q^{\prime},\varepsilon^{\prime},c^{\prime})$, where
@@
The condition (3.31) imposes a smallness constraint on the model parameters, a requirement that is commonly found in the literature. For instance, the convergence of FP to the open-loop NE for linear-quadratic games on complete graphs requires that $T$ or $(c,q,\varepsilon-q^{2})$ are sufficiently small, or $a$ is sufficiently large [Hu21]. The convergence of the Deep BSDE method [EHJ17] for stochastic control problems with Lipschitz models requires a small time duration or small Lipschitz coefficients [HHL22], as well as the well-posedness of forward-backward stochastic differential equations (FBSDEs) [MWZZ15]. Condition (3.31) aligns with these requirements, and it is certainly not the strictest constraint and can be easily extended. As long as reasonable bounds are imposed on model parameters such that the function $h$ in equation (3.23) has a fixed point in $\mathbb{R}_{+}$, Theorem 3.7 and Corollary 3.8 remain valid and provide a standard way to find a condition, under which FP is guaranteed to converge. Lastly, we want to emphasize two points: (i) FP does not necessarily have to start from $F^{i,0}\equiv 0$. From the proof of Theorem 3.7, the convergence results still hold as long as $\sup_{t\in[0,T]}\|\sum_{j=1}^{N}F^{j,0}_{t}\|\leq x^{*}$. (ii) Condition (3.31) does not impose any restrictions on the number of players $N$. The convergence result still holds even when there is a large number of players.
@@
We are particularly interested in the connection between the graph structure and the convergence of fictitious play (FP). This relationship can be analyzed through $C_{L}$ defined in condition (3.31). Fixing the other model parameters, an increase in $\|L\|$ leads to an increase in $C_{L}$, making it more likely to violate condition (3.31). Taking a closer look, we note that any graph Laplacian $L$ must satisfy $0\leq\|L\|\leq 2$. If a graph has no isolated vertices, its graph Laplacian satisfies $\|L\|\geq\frac{N}{N-1}$ [Chu97]. A connected bipartite graph has the norm of the graph Laplacian achieving the upper bound $\|L\|=2$, while a complete graph achieves the lower bound $\|L\|=\frac{N}{N-1}$. From the definition of $g_{k}(t)$ and the contraction (3.30) in the proof of Theorem 3.7, we derive an upper bound for the convergence rate of FP:
$\sup_{t\in[0,T]}\sup_{i\in[N]}\|\Delta F^{i,k}_{t}\|\leq 2^{-k}e^{6T[(a+q)\|L\|+x^{*}+Nx^{*}]}\sup_{t\in[0,T]}\sup_{i\in[N]}\|\Delta F^{i,0}_{t}\|,\ \forall k\in\mathbb{N},$ (3.35)
which admits an exponential rate of convergence. Although the precise convergence rate of FP remains difficult to quantify, numerical experiments show a connection between the graph structure and the convergence rate of FP. We refer the readers to Section 5.2 for a more detailed discussion.
Despite the existence of more efficient numerical algorithms for solving the Nash equilibrium in the LQ case, FP provides a novel approach that explicitly constructs a solution to the coupled Riccati system (2.14) and establishes its existence interval, as shown in Corollary 3.11.
Corollary 3.11.
If condition (3.31) holds, then the coupled Riccati system (2.14) admits a unique solution on $[0,T]$.
Proof.
Fictitious play provides an explicit construction of the solution, as demonstrated in the proof of Theorem 3.7. By [Luk71, Theorem 2.3 and Section 2.4], the solution to the coupled Riccati system (2.14) is unique whenever it exists. Combining these results concludes the proof. ∎
To the best of our knowledge, the most general existence interval for the solution to the coupled Riccati system (2.14) is given in [PC79], where the range of $T$ explicitly depends on the number of players $N$. In contrast, our existence condition (3.31) holds for any $N\in\mathbb{N}$, even when $N$ is large. This highlights the advantages of our approach, which establishes the local well-posedness of coupled Riccati systems through the lens of fictitious play.